40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3874     Entry time: Sat Nov 6 00:49:04 2010
Author: Koji 
Type: Update 
Category: IOO 
Subject: IMC table work 

[Suresh, Koji]

- Removed MCT optics in the IMC chamber

- Rotated MC1 and MC3 in clock-wise to debias the YAW bias offsets (-5V and -8V to -1.5V and -0.5V).

- Adjusted insertion of the MC1 OSEMs so as to have the outputs of about 1.0V.

- Locked to TEM00. Trying to get the beams at the center of the mirror using Yuta's A2L.

ELOG V3.1.3-